OpenStudy (anonymous):

if the concentration of reactants are expressed in mol/L and time is expressed in seconds, what are the units for the rate constant K?

OpenStudy (anonymous):

What is the equation for the rate constant?

OpenStudy (anonymous):

rate = k[a][b]

OpenStudy (anonymous):

with t on the end

OpenStudy (anonymous):

We want to rate to have units of (mole/L)*s right?

OpenStudy (anonymous):

Sorry, we want units of mole/L-s

OpenStudy (anonymous):

Therefore, the rate constant must have units of \(1/s\)

OpenStudy (xishem):

Assuming the two reactants have reaction orders of 1, the rate constant needs to have units of: \[\frac{M}{s}=k*M*M\]\[\frac{M}{s}=\frac{1}{M*s}*M*M\]\[\frac{M}{s}=\frac{M}{s}\]Therefore, the rate constant must have units of: \[M^{-1}s^{-1}\]

OpenStudy (xishem):

Or, alternatively \[L·mol^{-1}s^{-1}\]